OpenStudy (anonymous):

Simplify: HELP PLZ 21 - 3 × 6 + 52 37 133 28 64

OpenStudy (michele_laino):

we have to do the multiplication first

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

what is 3*6=...?

OpenStudy (anonymous):

18

OpenStudy (michele_laino):

ok! now: what is 21-18=...?

OpenStudy (anonymous):

21-3x6+52 21-18+52 3+52= 55

OpenStudy (anonymous):

rekt

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

finally: what is 3+52=...?

OpenStudy (anonymous):

55

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

there is no 55

OpenStudy (michele_laino):

I see, nevertheless my procedure is right!

OpenStudy (anonymous):

huh?

OpenStudy (michele_laino):

please check the text of your exercise

OpenStudy (anonymous):

oh the 2 is more upper than the 5

OpenStudy (anonymous):

look at the top that not supposed to be 52

OpenStudy (michele_laino):

ok! then we have to do the power first, so what is: 5^2=...?

OpenStudy (anonymous):

idk

OpenStudy (michele_laino):

it is simple: 5^2=5*5=...?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

not 10 25

OpenStudy (michele_laino):

5*5=25

OpenStudy (anonymous):

5*5 = 25

OpenStudy (anonymous):

hey brah michele_laino you want to help me with my stats problem you human calculator ;)

OpenStudy (michele_laino):

next we have to do the multiplication: 3*6=18 then we have to do this: 21-18=3 finally: 3+5^2=3+25=...?

OpenStudy (michele_laino):

what is 3+25=...?

OpenStudy (anonymous):

28

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

how do i get this guy to help me^?

OpenStudy (leoh1998):

what do you need help with?

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 4 ÷ 4 + 5 × 3 - 5 13 16 11 12

OpenStudy (anonymous):

division comes first them multiplication then addition and subtraction

OpenStudy (michele_laino):

here, you have to do the multiplication and division first, so: what is: 4:4=...? 5*3=...?

OpenStudy (anonymous):

thanks the answer is 11

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 20 - 8 + 5 - 9 ÷ 3 14 20 12 0

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

divide first then just - + - again ur answer should be 14.

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 49 - 2 × (2 + 4) 282 49 37 98

OpenStudy (anonymous):

parenthesis first =6 then multiply by 2= 12 then 49-12 = 37

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 30 - (4 + 1) - 4 ÷ 2 23 25 14 18

OpenStudy (anonymous):

hello

OpenStudy (leoh1998):

23

OpenStudy (anonymous):

se the correct order of operations to solve the problem below. 44 - 4 × (4 + 3) 40 280 16 163

OpenStudy (leoh1998):

16

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 5 + 6 × 1 10 11 12 43

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 72 ÷ 4 × (11 - 8) 108 54 18 6

OpenStudy (leoh1998):

54

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 7 - 2 × 2 + 3 13 -3 6 14

OpenStudy (leoh1998):

6

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 6 + 6 ÷ 3 4 8 13 22

OpenStudy (leoh1998):

8

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 3 × 3 + 3 ÷ 3 - (5 + 3) 8 4 16 2

OpenStudy (leoh1998):

2

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 12 + 1 × 7 92 18 19 91

OpenStudy (leoh1998):

19

OpenStudy (anonymous):

Divide 20 to the 9th power by 5; then add 66. Which of the following expressions matches the sentence above? Divide 20 to the 9th power by 5; then add 66. Which of the following expressions matches the sentence above?

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i gave you a medal

OpenStudy (leoh1998):

what?

OpenStudy (anonymous):

Directions: Select all the correct answers. Read the statement, and identify the expressions that are equivalent. the sum of a number times 3 and 15 15 × n + 3 (n × 3) + 15 15 + 3n 3 × 15n (n + 15) × 3 15 + 3 × n

OpenStudy (leoh1998):

A B E

OpenStudy (anonymous):

thanks thats the last question so buy

OpenStudy (leoh1998):

bye

OpenStudy (anonymous):

Simplify: 4 × 32 + 16 ÷ 4 - (6 + 2) 36 32 24 40

OpenStudy (shaik0124):

A common technique for remembering the order of operations is the abbreviation "PEMDAS", which is turned into the phrase "Please Excuse My Dear Aunt Sally". It stands for "Parentheses, Exponents, Multiplication and Division, and Addition and Subtraction".

OpenStudy (anonymous):

i have that open on google

OpenStudy (shaik0124):

solve according to pedmas

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

can you just help me

OpenStudy (shaik0124):

so now solve accordingly u will get the answer

OpenStudy (shaik0124):

yeeah sure

OpenStudy (shaik0124):

4 × 32 + 16 ÷ 4 - (6 + 2) according to pedmas first solve parenthesis in this question parenthesis part is 6+2=8 now 4*32+16/4-8

OpenStudy (shaik0124):

in pedmas we done the p part and consider the next one e which stands for expoents.since there is no exponent part we move to the next on in pedmas i.e, division in this 4*32+16/4-8 the division part is 16/4 =4 4*32+4-8

OpenStudy (anonymous):

there is an exponent the 2 is

OpenStudy (shaik0124):

now comes the multiplication in pedmas 128+4-8

OpenStudy (anonymous):

3*3=9

OpenStudy (shaik0124):

2 is inside the parenthesis so it cant be exponent

OpenStudy (shaik0124):

132-8

OpenStudy (anonymous):

the one you think is 3^2

OpenStudy (shaik0124):

124

OpenStudy (shaik0124):

no 3*3 in this * -> multiplication

OpenStudy (shaik0124):

4 × 32 + 16 ÷ 4 - (6 + 2) parenthesis p 4 *32+16/4-8 division d 4*32+4-8 multiplication m 128+4-8 addition a 132-8 subtraction s 124

OpenStudy (anonymous):

the answer is 32

OpenStudy (shaik0124):

then there is something wrong in ur question please recheck

OpenStudy (anonymous):

that 32 in the problem is 3^2

OpenStudy (shaik0124):

ok

OpenStudy (shaik0124):

so u should clearly mention the question now u solve the exponent part after parenthesis part

OpenStudy (shaik0124):

done?

OpenStudy (shaik0124):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

your muslim shaik0124

OpenStudy (anonymous):

yeah im muslim i just now realized you have arabic on your pic